# Rotating axionic AdS4 black hole dressed with a scalar field

### v1

@@
Abstract
This paper presents a new four-dimensional axionic-charged rotating black hole with a scalar field, which is defined by a structural function coupling the axionic field and a scalar potential. This configuration is characterized by an integration constant and two parameters. The thermodynamic parameters are obtained via the Euclidean procedure, where the validity of the first law of thermodynamics is ensured. With the above information, we can see that this rotating configuration is a good laboratory from the perspective of the holographic superconductor, in which the angular parameter plays an important role.
@@ I Introduction
Scalar fields non-minimally coupled to gravity have been exhaustively studied in the context of black holes (BHs) for nearly fifty years since the seminal work of N. Bocharova, K. Bronnikov, V. Melnikov Bocharova:1970skc and J. Bekenstein Bekenstein:1974sf ; Bekenstein:1975ts , which connected it to the no-hair conjecture Israel:1967wq ; Israel:1967za ; Carter:1971zc ; Hawking:1971tu . Although for the asymptotically flat situation, the model has been subject to debate from a physical point of view, and even the existence of solutions Bronnikov:1978mx ; Xanthopoulos:1992fm ; Klimcik:1993cia , the inclusion of a cosmological constant allows us to explore static BH configurations with a scalar field with other asymptotic behaviors Martinez:2002ru ; Barlow:2005yd ; Martinez:1996gn ; Martinez:2005di ; Dotti:2007cp ; Charmousis:2009cm ; Anabalon:2012tu ; Bardoux:2012tr ; Ayon-Beato:2015ada ; Ayon-Beato:2024xgp ; Ayon-Beato:2023bzp ; Cisterna:2021xxq ; Bravo-Gaete:2021hza . For a scalar field represented by $\phi$, the Lagrangian $\mathcal{L}_{\phi}$ typically takes on the following structure:
@@
It is interesting to note that BHs with non-minimal coupling and angular momentum have become a nontrivial task due to the complexity of the field equations, which is reflected in the absence of rich literature about it (see Refs. Natsuume:1999at ; Correa:2012rc ; Astorino:2014mda ; Bardoux:2013swa ; Anabalon:2009qt ; Erices:2017izj ; Baake:2020tgk ; Cardenas:2022jtz ). Nevertheless, as shown in Zhao:2013isa ; Zou:2014gla , a spinning configuration can be derived for the three-dimensional scenario where the model is represented through the action
$\mathcal{S}[g_{\mu\nu},\phi]=\int d^{4}x\sqrt{-g}\left[\frac{1}{2}(R-2\Lambda)+\mathcal{L}_{\phi}\right],$ (2)
@@
$\mathcal{E}_{\mu\nu}$ ::=:= $G_{\mu\nu}+\Lambda g_{\mu\nu}-T^{\phi}_{\mu\nu}=0,$ (3)
$\mathcal{E}_{\phi}$ ::=:= $\square\phi-\xi R\phi-\frac{dU}{d\phi}=0,$ (4)
@@
Using this idea raises the natural question of exploring it for the four-dimensional scenario. This is, starting with a metric ansätze of the form
@@
where $t\in\mathbb{R}$, $r$ is non-negative, while we consider the scalar field as $\phi=\phi(r)$ and set $\xi=\frac{1}{6}$. Further, if we take into account the field equations, which are listed in the Appendix section, we can obtain the expression for the scalar field $\phi$ from the combinations $\mathcal{E}^{t}_{\phantom{a}t}-\mathcal{E}^{r}_{\phantom{a}r}=0=\mathcal{E}^{t}_{\phantom{a}\theta}$. Nevertheless, the combination $\mathcal{E}^{\theta}_{\phantom{a}\theta}-\mathcal{E}^{\varphi}_{\phantom{a}\varphi}=0$ yields
@@
fixing $N^{\theta}(r)=N_{0}$ as an integration constant and, through a coordinate transformation $\tilde{\theta}=\theta+N_{0}t$, the static configuration is naturally recovered.
@@
The exercise aforementioned demonstrates that the prior action proposal (2) is insufficient, thereby redirecting our efforts toward a suitable election for matter sources. One approach is to adopt the concepts from Bardoux:2012tr ; Bardoux:2012aw ; Caldarelli:2013gqa , where two 3-forms that originate from Kalb-Ramond potentials are introduced, allowing us to obtain a planar version of the hairy BH given in Martinez:2002ru ; Martinez:2005di . In addition, it has recently been shown that when the source is composed of two axionic fields that vary linearly with the coordinates of the planar base manifold, charged planar Anti-de Sitter (AdS) BHs can arise Cisterna:2018hzf ; Cisterna:2019uek , enabling the computation of the DC conductivities due to the linear dependence on the coordinates for the axionic fields, providing a momentum dissipation mechanism Andrade:2013gsa ; Donos:2014cya ; Donos:2013eha , and even the potential to explore other contexts (see, for instance, Refs. Ling:2016lis ; Cisterna:2017jmv ; Wang:2018hwg ; Bhatnagar:2017twr ; BravoGaete:2019rci ; Hernandez-Vera:2024zui ). In particular, as we will see in the following lines, a four-dimensional rotating BHs configuration arises when for the action (2) we consider the inclusion of only one axionic field $\psi$ coupled with a function $\varepsilon(\phi)$:
$\mathcal{S}_{\psi}[g_{\mu\nu},\psi]=-\frac{1}{2}\int d^{4}x\sqrt{-g}\,{\varepsilon(\phi)}{\nabla_{\mu}\psi\nabla^{\mu}\psi}.$ (8)
@@
On the other hand, over the past few decades, the study of black holes in AdS spacetime has gained significant interest due to its crucial role in the AdS/Conformal Field Theory (CFT) correspondence Maldacena:1997re ; 2840490 . This duality unveils a fascinating relationship between gravitational theory and a strongly coupled quantum field theory that lives at the conformal boundary of the bulk gravitational spacetime. Thus, it provides a unique approach to studying the non-perturbative regime of quantum field theory systems, where traditional methods often fall short. Examples of these systems are prevalent in the domains of condensed matter physics, in which high-temperature superconductors figure as prototype systems on which duality can be applied. In this regard, the seminal work in holographic superconductivity successfully modeled a high-temperature superconductor using an Abelian-Higgs-like system in a non-back reacted planar Schwarzschild-AdS4 BH background Hartnoll:2008vx . This groundbreaking model elucidated the mechanism of holographic superconductivity by investigating the onset of the superconducting state driven by a $U(1)$ symmetry, breaking, and analyzing the linear response of the system to electric perturbations.
@@
Subsequent studies incorporated an external magnetic field to the superconducting system, enabling a holographic description of the London equation and the Abrikosov vortex lattice, which is characteristic of type-II superconductors Nakano:2008xc ; Hartnoll:2008kx ; Montull:2009fe ; Maeda:2009vf ; Donos:2020viz ; delaCruz-Lopez:2024chw . Notably, the dynamics of these holographic vortex lattices have been explored using both numerical and analytical techniques, uncovering the effect of the external magnetic field on the arrangement and density of vortices Xia:2021jzh ; Herrera-Mendoza:2022whz . Further realizations of holographic superconductors have considered non-commutative BH backgrounds delaCruz-Lopez:2024ync , and the effects of non-relativistic scaling symmetries using BHs with non-standard asymptotic behaviors Bu:2012zzb ; Lu:2013tza ; Natsuume:2018yrg ; Zhao:2013pva . Additionally, holographic superconductivity on isotropic and anisotropic rotating backgrounds has been studied in Herrera-Mendoza:2022whz ; Herrera-Mendoza:2024vfj ; Sonner:2009fk ; Lin:2014tza ; Srivastav:2019ixc , revealing that the BH rotation affects the properties of the superconducting state in such a way that the condensation of the scalar field is suppressed with increasing rotation. In Herrera-Mendoza:2024vfj , a connection was suggested between the rotation of a 5-dimensional anisotropic BH and the damping effects of quasiparticles in a superconducting material.
@@
Taking all of the above into consideration, in the present paper (i) we will show that a four-dimensional rotating configuration arises upon the inclusion of an axionic field, represented by the action (8), in addition to a structured form for the potential $U(\phi)$. We will see that only one integration constant and two parameters determine this configuration. Along with the previously mentioned, via the Euclidean procedure Gibbons:1976ue ; Regge:1974zd the thermodynamic parameters are obtained, where the validity of the first law of thermodynamics is ensured for this spinning hairy BH. With the above information, we will see that (ii) this rotating BH is a good laboratory from the point of view of the holographic superconductor, where the angular parameter plays a significant role.
@@
Two fundamental dimensions underpin the novelty of the current work. Firstly, the inclusion of an axionic field (8) in the toy model (2)
$\mathcal{S}_{\tiny{\mbox{total}}}[g_{\mu\nu},\phi,\psi]=\mathcal{S}+\mathcal{S}_{\psi},$ (9)
[…]
$-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-\frac{1}{2}\varepsilon(\phi)\nabla_{\mu}\psi\nabla^{\mu}\psi\equiv-\frac{1}{2}g_{ab}\nabla_{\mu}\Sigma^{a}\nabla^{\mu}\Sigma^{b},\hskip 8.53581pt\Sigma^{a}=(\phi,\psi),$
@@
which, in general, is not constant. Unlike the typical situation in supergravity models, where scalar manifolds are often realized as symmetric coset spaces with constant curvature Cremmer:1978km ; Andrianopoli:1996cm , the geometry emerging in this case does not correspond to any standard maximally symmetric space, or their noncompact counterparts Freedman:2012zz . This makes the existence of such a configuration particularly intriguing, suggesting a nontrivial interplay between the scalar sector and the underlying dynamics. Secondly, the inclusion of a rotational component offers an intriguing opportunity to explore holographic superconductors within the framework of the AdS/CFT correspondence.
@@ II Four-dimensional spinning axionic black holes dressed with a scalar field
Considering the total action (9) in four dimensions with $\xi=\frac{1}{6}$ and the equations of motion
$\mathcal{G}_{\mu\nu}$ ::=:= $\mathcal{E}_{\mu\nu}-T^{\psi}_{\mu\nu}=0,\qquad\mathcal{G}_{\phi}:=\mathcal{E}_{\phi}-\frac{d\varepsilon}{d\phi}{(\nabla_{\mu}\psi\nabla^{\mu}\psi)}=0,\qquad\nabla_{\mu}\left(\varepsilon(\phi)\nabla^{\mu}\psi\right)=0,$ (10)
with the corresponding contribution to the energy-momentum tensor
$T^{\psi}_{\mu\nu}=\varepsilon(\phi)\left(\nabla_{\mu}\psi\nabla_{\nu}\psi-\dfrac{1}{2}g_{\mu\nu}{\nabla_{\sigma}\psi\nabla^{\sigma}\psi}\right),$ (11)
the solution for the metric (6) is given by
@@
${N^{\theta}(r)}$ $=$ ${\frac{\sqrt{\alpha}}{192}\left[\ln\left(\dfrac{2B+r}{r}\right)+\dfrac{2B(r+B)(2B^{2}+2Br-r^{2})}{r^{4}}\right]},$ (13)
together with
[…]
and a negative cosmological constant
$\Lambda=-\frac{3}{\ell^{2}}.$ (15)
Moreover, the configuration is supported by a self-interacting potential $U(\phi)$ and coupling function $\varepsilon(\phi)$ which read:
@@
Here, $\eta>0$, $\alpha>0$ are coupling constants, while $B>0$ is the only integration constant present in the solution.
Several key observations can be made regarding the previously presented spinning solution. Firstly, we note that the rotating parameter $\alpha$ takes an important role at the moment to explore this configuration. In fact, at the limit $\alpha\rightarrow 0^{+}$ with $\eta\neq 0$, from equations (13) and (17), the rotating function $N^{\theta}$ as well as the coupling function $\varepsilon$ become zero, while that the potential takes the form (16) with $U_{\alpha}(\phi)=0$. Additionally, we note that the metric function $f$ and the scalar field $\phi$ do not vanish, allowing us to obtain a static configuration where the line element is given by
@@
Here, $T^{\phi}_{\mu\nu}$ and $T^{\psi}_{\mu\nu}$ were given previously in eqs. (5) and (11) respectively. It is interesting to note that eq. (23) have brought a lot of attention during recent years, allowing you to explore new and interesting solutions (see for example Refs. AyonBeato:2004ig ; AyonBeato:2005tu ; Ayon-Beato:2015qfa ; Ayon-Beato:2013bsa ; BravoGaete:2013djh ; Hassaine:2013cma ; Bernardo:2019yxp ; Hassaine:2006gz ; Alvarez:2016qky ; Erices:2024iah ).
With respect to the scalar field $\phi$, as indicated in equation (14), we note that tends to zero for large $r$. Additionally, when we perform an expansion for the coupling function $\varepsilon(\phi)$ and the potential $U(\phi)$, given in eqs. (16)-(19), we obtain
@@
implying that these expressions vanish trivially when $r\rightarrow+\infty$. In relation to the metric function $f$ and $N^{\theta}$ as defined in eqs. (12)-(13), it is worth noting that at infinity
@@
Here, for our notations (${}^{\prime})$ denotes the derivative with respect to coordinate $r$. Then, at the moment to evaluate (27) at $f$ and $N^{\theta}$, we get that around $r=0$: [^1] It is worth mentioning that there exists the situation where $r=-2B<0$. However, this case is dismissed due to the constraints imposed by the range of the radial coordinate $r$.

### v2

# A new rotating axionic AdS4 black hole dressed with a scalar field

@@
A new rotating axionic AdS4 black hole dressed with a scalar field
Moisés Bravo-Gaete1 moisesbravog@gmail.com, mbravo@ucm.cl    Fabiano F. Santos2,3 fabiano.ffs23@gmail.com    Jhony A. Herrera-Mendoza4,5 jhonyahm@gmail.com    Daniel F. Higuita-Borja5,6 dfhiguit@gmail.com 1Departamento de Matemática, Física y Estadística, Facultad de Ciencias Básicas, Universidad Católica del Maule, Casilla 617, Talca, Chile. 2School of Physics, Damghan University, Damghan, 3671641167, Iran. 3 Departamento de Física, Universidade Federal do Maranhão, São Luís, 65080-805, Brazil. 4Escuela de Física, Facultad de Ciencias, Universidad Nacional Autónoma de Honduras, Blvr. Suyapa, Tegucigalpa, Municipio del Distrito Central 11101, Honduras. 5Instituto de Física, Benemérita Universidad Autónoma de Puebla, Edificio IF-1, Ciudad Universitaria, Puebla, Pue. 72570, México, 6Instituto de Física, Universidad de Antioquia, Calle 70 No 52-21, Medellín, Colombia.
Resumo
This paper presents a new four-dimensional axionically charged rotating black hole with a scalar field, which is defined by a structural function coupling the axionic field and a scalar potential. This configuration is characterized by an integration constant and two constant parameters. The thermodynamic quantities are obtained via the Euclidean procedure, where the validity of the first law of thermodynamics is ensured. These results indicate that the rotating configuration provides a useful framework for exploring holographic superconductors, where the angular constant parameter plays a central role.
@@ I Introduction
Scalar fields non-minimally coupled to gravity have been a central topic in black hole (BH) physics since the pioneering works of N. Bocharova, K. Bronnikov, V. Melnikov Bocharova:1970skc and J. Bekenstein Bekenstein:1974sf ; Bekenstein:1975ts , which connected such configurations to the no-hair conjecture Israel:1967wq ; Israel:1967za ; Carter:1971zc ; Hawking:1971tu . Although for the asymptotically flat situation the model has been subject to debate from a physical point of view, and even the existence of solutions Bronnikov:1978mx ; Xanthopoulos:1992fm ; Klimcik:1993cia , the inclusion of a cosmological constant allows us to explore static BH configurations with a scalar field with other asymptotic behaviors Martinez:2002ru ; Barlow:2005yd ; Martinez:1996gn ; Martinez:2005di ; Dotti:2007cp ; Charmousis:2009cm ; Anabalon:2012tu ; Bardoux:2012tr ; Ayon-Beato:2015ada ; Ayon-Beato:2024xgp ; Ayon-Beato:2023bzp ; Cisterna:2021xxq ; Bravo-Gaete:2021hza . For a scalar field represented by $\phi$, the Lagrangian $\mathcal{L}_{\phi}$ typically takes on the following structure:
@@
The construction of BHs with non-minimal scalar couplings and rotation is technically challenging, which explains the scarcity of exact solutions in the literature (see Refs. Natsuume:1999at ; Correa:2012rc ; Astorino:2014mda ; Bardoux:2013swa ; Anabalon:2009qt ; Erices:2017izj ; Baake:2020tgk ; Cardenas:2022jtz ; Barrientos:2022yoz ; Bravo-Gaete:2025xya ). Nevertheless, as shown in Zhao:2013isa ; Zou:2014gla , a spinning configuration can be derived for the three-dimensional scenario where the model is represented through the action
$\int d^{3}x\sqrt{-g}\left[\frac{1}{2}(R-2\Lambda)+\mathcal{L}_{\phi}\right]=\int d^{3}x\sqrt{-g}(\mathcal{L}_{g}+\mathcal{L}_{\phi}),$ (2)
@@
$\mathcal{E}_{\mu\nu}$ $:=$ $G_{\mu\nu}+\Lambda g_{\mu\nu}-T^{\phi}_{\mu\nu}=0,$ (3)
$\mathcal{E}_{\phi}$ $:=$ $\square\phi-\xi R\phi-\frac{dU}{d\phi}=0,$ (4)
@@
This motivates the search for analogous rotating configurations in four dimensions; that is, starting with a metric ansatz of the form Klemm:1997ea ; Lemos:1995cm
@@
where $t\in\mathbb{R}$, $r$ is non-negative, while we consider the scalar field as $\phi=\phi(r)$ and set $\xi=\frac{1}{6}$. Here we note that the line element (6) describes a configuration with a planar (or toroidal) horizon geometry, and therefore the metric component $g_{\varphi\varphi}$ does not include a factor of $\sin^{2}\theta$, as would be the case for the spherical situation. Further, if we take into account the field equations, which are listed in Appendix A, we can obtain the expression for the scalar field $\phi$ from the combinations $\mathcal{E}^{t}_{\phantom{a}t}-\mathcal{E}^{r}_{\phantom{a}r}=0=\mathcal{E}^{t}_{\phantom{a}\theta}$. Nevertheless, the combination $\mathcal{E}^{\theta}_{\phantom{a}\theta}-\mathcal{E}^{\varphi}_{\phantom{a}\varphi}=0$ yields
@@
fixing $N^{\theta}(r)=N_{0}$ as an integration constant and, through a coordinate transformation $\tilde{\theta}=\theta+N_{0}t$, the static configuration is naturally recovered. Here, for our notations, $(^{\prime})$ denotes the derivative with respect to coordinate $r$.
@@
This result shows that the previous action (2) is insufficient, motivating us to search for more appropriate matter sources. One approach is to adopt the concepts from Bardoux:2012tr ; Bardoux:2012aw ; Caldarelli:2013gqa , where two 3-forms that originate from Kalb-Ramond potentials are introduced, allowing us to obtain a planar version of the hairy BH given in Martinez:2002ru ; Martinez:2005di . In addition, it has recently been shown that when the source is composed of two axionic fields that vary linearly with the coordinates of the planar base manifold, charged planar Anti-de Sitter (AdS) BHs can arise Cisterna:2018hzf ; Cisterna:2019uek , enabling the computation of the DC conductivities due to the linear dependence on the coordinates for the axionic fields, providing a momentum dissipation mechanism Andrade:2013gsa ; Donos:2014cya ; Donos:2013eha , and even the potential to explore other contexts (see, for instance, Refs. Ling:2016lis ; Cisterna:2017jmv ; Wang:2018hwg ; Bhatnagar:2017twr ; BravoGaete:2019rci ; Hernandez-Vera:2024zui ). The addition of a scalar field with an axionic profile has allowed the study of black strings in dynamical Chern-Simons modified gravity and their thermodynamics Corral:2021tww ; Cisterna:2018jsx , as well as in more general toy models, such as Horndeski Arratia:2020hoy . In this work, we show that a new four-dimensional rotating configuration arises when the action (2) is supplemented by a single axionic field $\psi$ coupled via a function $\varepsilon(\phi)$:
$\mathcal{S}_{\psi}[g_{\mu\nu},\psi]=-\frac{1}{2}\int d^{4}x\sqrt{-g}\,{\varepsilon(\phi)}{\nabla_{\mu}\psi\nabla^{\mu}\psi},$ (8)
where the term axionic refers to a shift-symmetric scalar field $\psi\rightarrow\psi+\mathrm{const.}$ that varies linearly with a spatial coordinate. This terminology was introduced in Andrade:2013gsa ; Donos:2014cya ; Donos:2013eha .
@@
On the other hand, over the past few decades, the study of BHs in AdS spacetime has gained significant interest due to its crucial role in the AdS/Conformal Field Theory (CFT) correspondence Maldacena:1997re ; 2840490 . This duality unveils a fascinating relationship between a gravitational theory and a strongly coupled quantum field theory that lives at the conformal boundary of the bulk gravitational spacetime. This duality provides a framework to explore strongly coupled quantum systems beyond perturbative methods. Examples of these systems are prevalent in the domains of condensed matter physics, in which high-temperature superconductors figure as prototype systems on which duality can be applied. In this regard, the seminal work in holographic superconductivity successfully modeled a high-temperature superconductor using an Abelian-Higgs-like system in a non-back-reacted planar Schwarzschild-AdS4 BH background Hartnoll:2008vx . This groundbreaking model elucidated the mechanism of holographic superconductivity by investigating the onset of the superconducting state driven by a $U(1)$ symmetry breaking and analyzing the linear response of the system to electric perturbations.
@@
Subsequent studies incorporated an external magnetic field into the superconducting system, enabling a holographic description of the London equation and the Abrikosov vortex lattice, which is characteristic of type-II superconductors Nakano:2008xc ; Hartnoll:2008kx ; Montull:2009fe ; Domenech:2010nf ; Maeda:2009vf ; Montull:2011im ; Montull:2012fy ; Salvio:2012at ; Donos:2020viz ; delaCruz-Lopez:2024chw . Notably, the dynamics of these holographic vortex lattices have been explored using both numerical and analytical techniques, uncovering the effect of the external magnetic field on the arrangement and density of vortices Xia:2021jzh ; Herrera-Mendoza:2022whz . Further realizations of holographic superconductors have considered non-commutative BH backgrounds delaCruz-Lopez:2024ync , and the effects of non-relativistic scaling symmetries using BHs with non-standard asymptotic behaviors Bu:2012zzb ; Lu:2013tza ; Natsuume:2018yrg ; Zhao:2013pva . Additionally, holographic superconductivity on isotropic and anisotropic rotating backgrounds has been studied in Herrera-Mendoza:2022whz ; Herrera-Mendoza:2024vfj ; Sonner:2009fk ; Lin:2014tza ; Srivastav:2019ixc , revealing that the BH rotation affects the properties of the superconducting state in such a way that the condensation of the scalar field is suppressed with increasing rotation. In Herrera-Mendoza:2024vfj , a connection was suggested between the rotation of a 5-dimensional anisotropic BH and the damping effects of quasiparticles in a superconducting material.
@@
Taking all of the above into consideration, in the present paper, we will show that (i) a new four-dimensional rotating configuration arises upon the inclusion of an axionic field, represented by the action (8), in addition to a structured form for the potential $U(\phi)$. In our construction, the scalar potential and the coupling function $\varepsilon(\phi)$ were not postulated a priori but reconstructed consistently to ensure integrability of the system, playing a key role in supporting the rotating configuration and exhibiting nontrivial behavior. This reconstruction method has been performed in the literature (e.g., Erices:2017izj ; Cisterna:2019uek ; Cisterna:2018hzf )). The resulting BH depends only on an integration constant and two constant parameters. Along with the previously mentioned, via the Euclidean procedure Gibbons:1976ue ; Regge:1974zd , the thermodynamic quantities are obtained, where the validity of the first law of thermodynamics is ensured for this spinning hairy BH. With the above information, we will see that (ii) this rotating BH is a good laboratory for a holographic perspective, where the condensation and electrical conductivity represent new behavior associated with the interaction of rotation, momentum dissipation, and non-minimal couplings, revealing effects originating from the rotating axionic background.
@@
Two fundamental dimensions underpin the novelty of the current work. Firstly, the inclusion of an axionic field (8) in the toy model (2):
$\mathcal{S}_{\tiny{\mbox{total}}}[g_{\mu\nu},\phi,\psi]=\int d^{4}x\sqrt{-g}(\mathcal{L}_{g}+\mathcal{L}_{\phi})+\mathcal{S}_{\psi},$ (9)
provides opportunities to obtain a rotating configuration as well as allows us to study the thermodynamic properties outside the three-dimensional scenario. It is noteworthy that the kinematic sector of the Lagrangian density can be written as a nonlinear $\sigma$-model Gell-Mann:1960mvl
$\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-\frac{1}{2}\varepsilon(\phi)\nabla_{\mu}\psi\nabla^{\mu}\psi\equiv-\frac{1}{2}g_{ab}\nabla_{\mu}\Sigma^{a}\nabla^{\mu}\Sigma^{b},\hskip 8.53581pt\Sigma^{a}=(\phi,\psi),$
@@
which, in general, is not constant. Unlike the typical situation in supergravity models, where scalar manifolds are often realized as symmetric coset spaces with constant curvature Cremmer:1978km ; Andrianopoli:1996cm , the geometry emerging in this case does not correspond to any standard maximally symmetric space or its non-compact counterparts Freedman:2012zz . This makes the existence of such a configuration particularly intriguing, suggesting a nontrivial interplay between the scalar sector and the underlying dynamics. Secondly, the inclusion of a rotational component offers an intriguing opportunity to explore holographic superconductors within the framework of the AdS/CFT correspondence.
@@ II Four-dimensional spinning axionic black holes dressed with a scalar field
Consider the total action (9) in four dimensions with $\xi=\frac{1}{6}$. The equations of motion are given by
$\mathcal{G}_{\mu\nu}$ $:=$ $\mathcal{E}_{\mu\nu}-T^{\psi}_{\mu\nu}=0,\quad\mathcal{G}_{\phi}:=\mathcal{E}_{\phi}-\frac{d\varepsilon}{d\phi}{(\nabla_{\mu}\psi\nabla^{\mu}\psi)}=0,\quad\mathcal{G}_{\psi}=\nabla_{\mu}\left(\varepsilon(\phi)\nabla^{\mu}\psi\right)=0,$ (10)
where the corresponding contribution to the energy-momentum tensor takes the form:
$T^{\psi}_{\mu\nu}=\varepsilon(\phi)\left(\nabla_{\mu}\psi\nabla_{\nu}\psi-\dfrac{1}{2}g_{\mu\nu}{\nabla_{\sigma}\psi\nabla^{\sigma}\psi}\right).$ (11)
For completeness, the explicit expressions for $T^{\psi}_{\mu\nu}$ are provided in Appendix A. Firstly, following Refs. Cisterna:2018hzf ; Cisterna:2019uek ; Andrade:2013gsa ; Donos:2014cya ; Donos:2013eha , we consider an axionic field $\psi$ that varies linearly with respect to the coordinate $\varphi$, obtaining that $\mathcal{G}_{\psi}=\varepsilon(\partial^{2}_{\varphi}\psi/r^{2})=0$, where $\partial^{2}_{\varphi}\psi:=\partial^{2}\psi/\partial\varphi^{2}$, yielding:
$\psi(\varphi)=B\varphi$}. (12)
Here, $B>0$ is an integration constant. Then, considering the combinations of the Einstein equations:
$\mathcal{G}_{t}^{\theta}$ $=$ $-\frac{r}{12}\left[r(\phi^{2}-6)(N^{\theta})^{\prime\prime}+2(N^{\theta})^{\prime}(2\phi^{2}+r\phi\phi^{\prime}-12)\big)\right]=0,$ (13)
$\mathcal{G}_{t}^{t}-\mathcal{G}_{r}^{r}$ $=$ $-\frac{1}{12}N^{\theta}r^{2}(\phi^{2}-6)(N^{\theta})^{\prime\prime}-\frac{1}{3}f\phi\phi^{\prime\prime}-\frac{1}{6}rN^{\theta}(r\phi\phi^{\prime}+2\phi^{2}-12)(N^{\theta})^{\prime}$ (14)
$+$ $\frac{2}{3}(\phi^{\prime})^{2}f=0,$
allow us to obtain an explicit expression for the scalar field $\phi$, via the differential equation
$\frac{f}{3}\Big(\phi\phi^{\prime\prime}-2(\phi^{\prime})^{2}\Big)=0,$
given by [^1] Here we note that the most general solution is $\phi(r)=1/(ar+b)$, with $a$ and $b$ integration constants. Nonetheless, the remaining equations of motion demonstrate a relationship between these constants and the integration constant $B$.
$\phi(r)=\dfrac{\sqrt{6}B}{r+B}.$ (15)
From eq. (13), we obtain the expression for the rotating function $N^{\theta}$, which reads
${N^{\theta}(r)}$ $=$ ${\frac{\sqrt{\alpha}}{192}\left[\ln\left(\dfrac{2B+r}{r}\right)+\dfrac{2B(r+B)(2B^{2}+2Br-r^{2})}{r^{4}}\right]}$ (16)
$=$ $\frac{\sqrt{\alpha}}{192}\left[\ln\left(\frac{\sqrt{6}+\phi(r)}{\sqrt{6}-\phi(r)}\right)+\dfrac{2\sqrt{6}(4\sqrt{6}\phi(r)-\phi(r)^{2}-6)\phi(r)}{(\sqrt{6}-\phi(r))^{4}}\right],$
where $\alpha$ is a constant parameter, while the structural function is obtained via the combination $\mathcal{G}_{\theta}^{\theta}-\mathcal{G}_{\varphi}^{\varphi}=0,$ which is proportional to the algebraic equation:
$12(\phi+\sqrt{6})(\sqrt{6}-\phi)^{5}\varepsilon(\phi)-\alpha\phi^{4}=0,$
and taking the form:
$\varepsilon(\phi)$ $={\dfrac{\alpha\phi^{4}}{12(\sqrt{6}-\phi)^{5}(\sqrt{6}+\phi)}}.$ (17)
Finally, the solution for the metric (6) with $N^{\theta}$ given in eq. (16), through
$\mathcal{G}_{t}^{t}-\mathcal{G}_{\theta}^{\theta}$ $=$ $-72r^{8}(r+B)\left(B+\frac{r}{2}\right)^{2}f^{\prime\prime}-72r^{7}\left(B+\frac{r}{2}\right)B^{2}f^{\prime}$ (18)
$+$ $36r^{6}(r+2B)(r^{2}+3Br+4B^{2})f+\alpha B^{6}(r+B)^{5}=0,$
reads
@@
where there $\eta$ is another constant parameter and $\ell$ is the AdS radius. With a negative cosmological constant,
$\Lambda=-\frac{3}{\ell^{2}},$ (20)
the configuration is supported by a self-interacting potential $U(\phi)$ (this is $\mathcal{G}_{\mu}^{\nu}=0$ or $\mathcal{G}_{\phi}=0$), given by
@@
The coupling function $\varepsilon(\phi)$ (eq. (17)) and the potential $U(\phi)$ (eqs. (21) and (22)-(23)) play a providential role in supporting the rotating configuration (see eqs. (6), (19)-(20)). It is important to emphasize that in an analogous way to the three-dimensional situation Zhao:2013isa ; Zou:2014gla , within the ansätze (6) under the assumptions $\phi=\phi(r)$, $\psi=\psi(\varphi)$ together with $\xi=\frac{1}{6}$, the field equations (10)-(11) admit a novel rotating configuration described by the functions $\phi$, $\psi$, $N^{\theta}$, $U(\phi)$, $\varepsilon(\phi)$, and $f$, given by an only integration constant $B>0$ as well as the constant parameters $\alpha>0$ and $\eta>0$.
The obtained new spinning solution exhibits the following relevant physical properties. Firstly, we note that the rotating parameter $\alpha$ takes an important role at the moment to explore this configuration. In fact, at the limit $\alpha\rightarrow 0^{+}$ with $\eta\neq 0$, from equations (16) and (17), the rotating function $N^{\theta}$ as well as the coupling function $\varepsilon$ become zero, while the potential takes the form (21) with $U_{\alpha}(\phi)=0$. Additionally, we note that the metric function $f$ and the scalar field $\phi$ do not vanish, allowing us to obtain a new static configuration where the line element is given by
@@
Here, $T^{\phi}_{\mu\nu}$ and $T^{\psi}_{\mu\nu}$ were given previously in eqs. (5) and (11) respectively. It is important to note that eq. (27) have brought a lot of attention during recent years, allowing you to explore new and interesting solutions (see, for example, Refs. AyonBeato:2004ig ; AyonBeato:2005tu ; Ayon-Beato:2015qfa ; Ayon-Beato:2013bsa ; BravoGaete:2013djh ; Hassaine:2013cma ; Bernardo:2019yxp ; Hassaine:2006gz ; Alvarez:2016qky ; Erices:2024iah ).
With respect to the scalar field $\phi$, as indicated in equation (15), we note that it is real, finite, and a monotonically decreasing function, where $\phi<6$ for $r>0$ and $B>0$, whereas
$\phi(r)\rightarrow 0\quad\mbox{as}\quad r\rightarrow+\infty.$
As we will see below, since the horizon is located at $r=r_{h}>0$, the physically relevant region satisfies $0<\phi<\sqrt{6}$, and the apparent pole of the functions $\varepsilon(\phi)$ and $U(\phi)$ at $\phi^{2}=6$ lies at $r=r_{s}=0$, which corresponds to the curvature singularity inside the spinning BH. Additionally, when we perform an expansion for the coupling function $\varepsilon(\phi)$ and the potential $U(\phi)$, given in eqs. (17) and (21)-(23), we obtain:
@@
implying that these expressions vanish trivially when $r\rightarrow+\infty$. In relation to the metric function $f$ and $N^{\theta}$ as defined in eqs. (16) and (19), it is worth noting that at infinity:
@@
Then, at the moment to evaluate (31) at $f$ and $N^{\theta}$, we get that around r=0r=0: [^2] It is worth mentioning that there exists the situation where r=2B<0r=-2B<0. However, this case is dismissed due to the constraints imposed by the range of the radial coordinate $r$.